OpenStudy (anonymous):

Help please ill give medal!!!! A function rule C= 10n +26 relates the number of people n who attended a small concert to the cost in dollars C of the concert. make a table of the input/output pairs to show the cost of 27, 39, and 43 people attended

OpenStudy (anonymous):

ill try ok

OpenStudy (anonymous):

For each of the values, substitute for n in 10n + 26: 10(27) + 26 = 10*27 + 26 = 270 + 26 = 296 10(39) + 26 = 10*39 + 26 = 390 + 26 = 416 10(43) + 26 = 10*43 + 26 = 430 + 26 = 456

OpenStudy (anonymous):

So i guess the input/output pairs are: 27/296 39/416 43/456

OpenStudy (anonymous):

Thank you so much!